According to the National Oceanic and Atmospheric Administration, the state of Michigan has an average of 18 tornadoes per year. Michigans land area is approximately 58,110 square miles. Find the probability that a randomly selected 1000 square mile region in Michigan will have exactly 2 tornadoes this year.
Answer #1

expected number of tornadoes in 1000 square mile =1000*18/58110=0.31

therefore from poisson approximation

probability that a randomly selected 1000 square mile region in Michigan will have exactly 2 tornadoes

=e-0.31*0.312/2! =0.0352 ( please try 0.0333 if expected value is to be taken with single decimal place)